# Linear complementary dual quasi-cyclic codes of index 2

### v1

@@ Abstract
We provide a polynomial approach to investigate linear complementary dual (LCD) quasi-cyclic codes over finite fields. We establish necessary and sufficient conditions for LCD quasi-cyclic codes of index 2 with respect to the Euclidean, Hermitian, and symplectic inner products. As a consequence of these characterizations, we derive necessary and sufficient conditions for LCD one-generator quasi-cyclic codes.
@@ 1 Introduction
In the 2000s, Ling and Solé studied the algebraic structure of quasi-cyclic codes in a series of articles [19, 20, 21, 22]. In [18], Lally and Fitzpatrick proved that every quasi-cyclic code has a generating set of polynomials in the form of a reduced Gröbner basis. Based on these structural properties, more asymptotic results, minimum distance bounds, and further applications of quasi-cyclic codes were obtained in the literature. To name a few, we refer to the paper Semenov and Trifonov [31] on the spectral method for quasi-cyclic codes, see also by other authors in [23] and [34]. One-generator quasi-cyclic codes were studied in [1], [27] and [28]. Applications of quasi-cyclic codes in constructing quantum codes have become a very active research topic in recent years, see for example [4], [9], and [12].
Linear codes with complementary duals (LCD codes) were introduced by Massey in [25]. In [29], Sendrier proved that LCD codes are asymptotically good and used them in relation to equivalence testing of linear codes in [30]. Recently, LCD codes became an attractive research interest as they offer solutions to many cryptographic problems, for example against side-channel attacks and fault non-invasive attacks, see [6]. In [7], it was shown that any linear code over $\mathbb{F}_{q}$ ($q>3$) is equivalent to a Euclidean LCD code and any linear code over $\mathbb{F}_{q^{2}}$ ($q>2$) is equivalent to a Hermitian LCD code.
In 1994, a characterization for LCD cyclic codes in terms of their generator polynomials was provided by Yang and Massey in [33]. For the case of quasi-cyclic codes, Esmaeili and Yari [8] provided a sufficient condition for quasi-cyclic codes to be Euclidean LCD codes and gave a method for constructing quasi-cyclic Euclidean LCD codes. In 2016, Güneri, Özkaya and Solé in [14] characterized Euclidean LCD quasi-cyclic codes using the Chinese Remainder Theorem (CRT) decomposition of codes introduced by Ling and Solé in [19]. Recently in [5] and [11], characterizations of Euclidean LCD one generator quasi-cyclic codes of index $\ell$ were obtained.
Based on the previous results of [14], [18], and [19], in this paper we provide a new characterization for LCD quasi-cyclic codes of index 2 in terms of generating sets of polynomials.
@@ 2.2 Decomposition of quasi-cyclic codes
For each $i$ and $j$, let $F_{i}=F[x]/(f_{i})$, $H_{j}^{\prime}=F[x]/(h_{j})$, and $H_{j}^{\prime\prime}=F[x]/(h_{j}^{*})$. Let $\xi$ be a primitive $m^{\text{th}}$ root of unity over $F$. Let $\xi^{u_{i}}$ and $\xi^{v_{j}}$ be roots of $f_{i}(x)$ and $h_{j}(x)$, respectively. Then we also have $h_{j}^{*}(\xi^{-v_{j}})=0$, and $F_{i}\cong F(\xi^{u_{i}})$, $H_{j}^{\prime}\cong F(\xi^{v_{j}})$, and $H_{j}^{\prime\prime}\cong F(\xi^{-v_{j}})$, see [15, p. 136].
The map ¯$\bar{\ }:f(x)\mapsto\bar{f}(x)$ can be naturally extended to the following isomorphisms:
¯$\begin{split}\bar{\ }:\ \mathbb{F}_{q}[x]/(f_{i}(x))&\to\mathbb{F}_{q}[x]/(f_{i}(x)),\\ \bar{\ }:\ \mathbb{F}_{q}[x]/(h_{j}(x))&\to\mathbb{F}_{q}[x]/(h_{j}^{*}(x)).\\ \end{split}$ (1)
@@
where $\textbf{c}=(c_{1}(\xi^{u_{i}}),\dots,c_{\ell}(\xi^{u_{i}})),\textbf{d}=(d_{1}(\xi^{u_{i}}),\dots,d_{\ell}(\xi^{u_{i}}))\in F_{i}^{\ell}.$ This is the inner product induced by $x\mapsto x^{-1}$, not the usual Hermitian inner product, see [21, p. 2693]. For each $1\leq j\leq p$, the vector space $(H_{j}^{\prime})^{\ell}\cong(H_{j}^{\prime\prime})^{\ell}$ is equipped with the usual Euclidean inner product and ${\perp_{e}}$ denotes the usual Euclidean dual.
@@
Since $f_{i}(x)$ is self-reciprocal, the cardinality of $F_{i}$, say $q_{i}$, is an even power of $q$ for all $1\leq i\leq s$ with two exceptions. One of these exceptions, for all $m$ and $q$, is the field coming from the irreducible factor $x-1$ of $x^{m}-1.$ When $q$ is odd and $m$ is even, $x+1$ is another self-reciprocal irreducible factor of $x^{m}-1$. In these cases, $q_{i}=q$. Except for two cases, the Hermitian inner product $\langle\cdot,\cdot\rangle_{h}$ is equivalent to the usual Hermitian inner product, see also [14, p.72] and [15, p.136]. For the two exceptions, in which case the corresponding field $F_{i}$ is $F$, we equip $F_{i}^{\ell}$ with the usual Euclidean inner product. Then the previous formula for $\langle\cdot,\cdot\rangle_{h}$ is still true, since $\xi^{u_{i}}=\xi^{-u_{i}}=\pm 1$.
We have the following characterization of Euclidean LCD QC codes from [15, Theorem 7.3.6].
@@ 2.3 Quasi-cyclic codes of index 2
Let $C$ be a quasi-cyclic code of length $2m$ and index $2$ Then $C$ is generated by two elements $(g_{11}(x),g_{12}(x))$ and $(0,g_{22}(x))$ such that they satisfy the following conditions:
@@
1. The code $C$ is generated by two elements $(g_{11}(x),g_{12}(x)\mod{g_{22}(x)})$ and $(0,g_{22}(x))$ satisfying Conditions ()($\ast$)( ∗ ).
@@
In [27, Lemma 1], Séguin showed that if $C$ is a quasi-cyclic code generated by one element $(g_{11}(x),g_{12}(x))$ with $g_{11}(x)\mid x^{m}-1$ and $g(x)=\gcd(g_{11}(x),g_{12}(x))$, then $\dim C=m-\deg g(x)$. With the choice of $g_{22}(x)$ described in Lemma 2.4, we see that the dimension of $C$ in Theorem 2.2 is consistent with the result by Séguin.
@@ 3 Euclidean LCD quasi-cyclic codes of index 2
Let $C$ be a quasi-cyclic code of index $2$. Let $(g_{11}(x),g_{12}(x))$ and $(0,g_{22}(x))$ be the generators of $C$ satisfying Conditions ()($\ast$)( ∗ ). Then $C$ is Euclidean LCD if and only if all of the following conditions are true:
@@
Assume that $l$ is not self-reciprocal. This implies that there exists $h_{j}$ such that $h_{j}\mid l$ and $h_{j}^{*}\nmid l$. Since $h_{j}\mid l$, it follows that $h_{j}\nmid\text{lcm}(g_{11},g_{22})$. In particular, $h_{j}\nmid g_{11}$ and $h_{j}\nmid g_{22}$. Then $g_{11}(\xi^{v_{j}})\neq 0$, $g_{22}(\xi^{v_{j}})\neq 0$ and ${\rm rank}(G_{j}^{\prime})=2$. This implies that $C_{j}^{\prime}=(H_{j}^{\prime})^{2}.$
@@ 1. Since $g$ and $l$ are self-reciprocal,
Furthermore, the reciprocal of $x^{m}-1$ is $-(x^{m}-1)$, so we can rewrite
@@
2. $h_{j}\mid l$. This condition implies that $h_{j}\nmid g$ and $h_{j}^{*}\nmid g$. Then $C_{j}^{\prime}$ and $C_{j}^{\prime\prime}$ are 2-dimensional, which implies that $C_{j}^{\prime\perp_{e}}$ and $C_{j}^{\prime\prime\perp_{e}}$ are 00-dimensional. Then $C_{j}^{\prime}\cap C_{j}^{\prime\prime\perp_{e}}=\{\mathbf{0}\}$ and $C_{j}^{\prime\prime}\cap C_{j}^{\prime\perp_{e}}=\{\mathbf{0}\}$.
@@ 4 One-generator quasi-cyclic codes of index 2
In view of Lemma 2.4, we can assume that $C$ is generated by two elements $(g_{11}(x),\tilde{g}_{12}(x))$ and $(0,g_{22}(x))$ satisfying Conditions ()($\ast$)( ∗ ), where $\tilde{g}_{12}(x)=g_{12}(x)\mod{g_{22}(x)}$.
@@ 2. The conditions $\gcd(g_{11}^{\prime},g_{12})=1$ and $\gcd(g_{11}^{\prime},\bar{g}_{12})=1$ imply that
We recall from Lemma 2.4 we can assume that the code $C$ is generated by two elements $(g_{11}(x),\tilde{g}_{12}(x))$ and $(0,g_{22}(x))$ satisfying Conditions ()($\ast$)( ∗ ), where
@@ 5 Symplectic LCD quasi-cyclic codes of index 2
Assume that all the notations are the same as in the previous sections. First, we recall some definitions and results, for details see [4, Section V]. For $x=(x_{1}|x_{2})$ and $y=(y_{1}|y_{2})$ in $F^{2m}$, where $x_{i},y_{i}\in F^{m}$ for $i=1,2$, we have
@@
Let $C$ be a QC code of length $2m$ and index $2$ with CRT decomposition given in (2). By extending the map $\tau$ canonically to the vector spaces $(F_{i})^{2},(H_{j}^{\prime})^{2},(H_{j}^{\prime\prime})^{2}$ and applying the maps component-wise to (3), we obtain symplectic dual of $C$ (see [4, Proposition V.1]).
@@
Using the above characterization, we have the following characterization of symplectic LCD QC codes in terms of constituents (see [4, Eq. V.7]).
@@
Let $C$ be a quasi-cyclic code of index $2$. Let $(g_{11}(x),g_{12}(x))$ and $(0,g_{22}(x))$ be the generators of $C$ satisfying Conditions ()($\ast$)( ∗ ). Then $C$ is symplectic LCD if and only if all of the following conditions are true:
@@
Furthermore, the reciprocal of $x^{m}-1$ is $-(x^{m}-1)$, so we can rewrite

### v2

@@ Abstract
We provide a polynomial approach to investigate linear complementary dual (LCD) quasi-cyclic codes over finite fields. We establish necessary and sufficient conditions for LCD quasi-cyclic codes of index 2 with respect to the Euclidean, Hermitian, and symplectic inner products. As a consequence of these characterizations, we derive necessary and sufficient conditions for LCD one-generator quasi-cyclic codes. Furthermore, using these characterizations, we construct some new quasi-cyclic LCD codes over small fields.
@@ 1 Introduction
In the 2000s, Ling and Solé studied the algebraic structure of quasi-cyclic codes in a series of articles [26, 27, 28, 29]. In [23], Lally and Fitzpatrick proved that every quasi-cyclic code has a generating set of polynomials in the form of a reduced Gröbner basis. Based on these structural properties, more asymptotic results, minimum distance bounds, and further applications of quasi-cyclic codes were obtained in the literature. To name a few, we refer to the paper Semenov and Trifonov [40] on the spectral method for quasi-cyclic codes, see also by other authors in [32] and [47]. Applications of quasi-cyclic codes in constructing quantum codes have become a very active research topic in recent years, see for example [11], [12], and [15].
Linear codes with complementary duals (LCD codes) were introduced by Massey in [34]. In [38], Sendrier proved that LCD codes are asymptotically good and used them in relation to equivalence testing of linear codes in [39]. Recently, LCD codes became an attractive research interest as they offer solutions to many cryptographic problems, for example against side-channel attacks and fault non-invasive attacks, see [6]. Several constructions of LCD codes with optimal parameters are known, see for example [3, 5, 7, 30, 41, 45]. In [8], it was shown that any linear code over $\mathbb{F}_{q}$ ($q>3$) is equivalent to a Euclidean LCD code and any linear code over $\mathbb{F}_{q^{2}}$ ($q>2$) is equivalent to a Hermitian LCD code. Analogous results were considered for symplectic self-dual codes in [25] and vector rank metric codes in [19].
In 1994, a characterization for LCD cyclic codes in terms of their generator polynomials was provided by Yang and Massey in [46]. For the case of quasi-cyclic codes, Esmaeili and Yari [10] provided a sufficient condition for quasi-cyclic codes to be Euclidean LCD codes and gave a method for constructing quasi-cyclic Euclidean LCD codes. In 2016, Güneri, Özkaya and Solé in [17] characterized Euclidean LCD quasi-cyclic codes using the Chinese Remainder Theorem (CRT) decomposition of codes introduced by Ling and Solé in [26].
One-generator quasi-cyclic codes were studied in [1], [36] and [37]. The general structure of quasi-cyclic codes of index 2 was examined in [2]. Recently in [4] and [14], characterizations of LCD one-generator quasi-cyclic codes of index $\ell$ were obtained.
Based on the previous results of [2], [17], [23], and [26], in this paper we provide a new characterization for Euclidean, Hermitian and symplectic LCD quasi-cyclic codes of index 2 in terms of generating sets of polynomials. Our results extend the existing characterizations obtained in [4] and [14] for one-generator quasi-cyclic codes of index 2.
@@ 2.2 Decomposition of quasi-cyclic codes
For each $i$ and $j$, let $F_{i}=F[x]/(f_{i})$, $H_{j}^{\prime}=F[x]/(h_{j})$, and $H_{j}^{\prime\prime}=F[x]/(h_{j}^{*})$. Let $\xi$ be a primitive $m^{\text{th}}$ root of unity over $F$. Let $\xi^{u_{i}}$ and $\xi^{v_{j}}$ be roots of $f_{i}(x)$ and $h_{j}(x)$, respectively. Then we also have $h_{j}^{*}(\xi^{-v_{j}})=0$, and $F_{i}\cong F(\xi^{u_{i}})$, $H_{j}^{\prime}\cong F(\xi^{v_{j}})$, and $H_{j}^{\prime\prime}\cong F(\xi^{-v_{j}})$, see [18, p. 136].
The map $\bar{\ }:f(x)\mapsto\bar{f}(x)$ can be naturally extended to the following isomorphisms:
[…]
@@
where $\textbf{c}=(c_{1}(\xi^{u_{i}}),\dots,c_{\ell}(\xi^{u_{i}})),\textbf{d}=(d_{1}(\xi^{u_{i}}),\dots,d_{\ell}(\xi^{u_{i}}))\in F_{i}^{\ell}.$ This is the inner product induced by $x\mapsto x^{-1}$, not the usual Hermitian inner product, see [28, p. 2693]. For each $1\leq j\leq p$, the vector space $(H_{j}^{\prime})^{\ell}\cong(H_{j}^{\prime\prime})^{\ell}$ is equipped with the usual Euclidean inner product and $\perp_{e}$} denotes the usual Euclidean dual.
@@
Since $f_{i}(x)$ is self-reciprocal, the cardinality of $F_{i}$, say $q_{i}$, is an even power of $q$ for all $1\leq i\leq s$ with two exceptions. One of these exceptions, for all $m$ and $q$, is the field coming from the irreducible factor x1x-1 of $x^{m}-1.$ When $q$ is odd and $m$ is even, $x+1$ is another self-reciprocal irreducible factor of $x^{m}-1$. In these cases, $q_{i}=q$. Except for two cases, the Hermitian inner product $\langle\cdot,\cdot\rangle_{h}$ is equivalent to the usual Hermitian inner product, see also [17, p.72] and [18, p.136]. For the two exceptions, in which case the corresponding field $F_{i}$ is $F$, we equip $F_{i}^{\ell}$ with the usual Euclidean inner product. Then the previous formula for $\langle\cdot,\cdot\rangle_{h}$ is still true, since $\xi^{u_{i}}=\xi^{-u_{i}}=\pm 1$.
We have the following characterization of Euclidean LCD QC codes from [18, Theorem 7.3.6].
@@ 2.3 Quasi-cyclic codes of index 2
Let $C$ be a quasi-cyclic code of length $2m$ and index $2.$ Then $C$ is generated by two elements $(g_{11}(x),g_{12}(x))$ and $(0,g_{22}(x))$ such that they satisfy the following conditions:
@@
1. The code $C$ is generated by two elements $(g_{11}(x),g_{12}(x)\mod{g_{22}(x)})$ and $(0,g_{22}(x))$ satisfying Conditions $(\ast)$.
@@
In [36, Lemma 1], Séguin showed that if $C$ is a quasi-cyclic code generated by one element $(g_{11}(x),g_{12}(x))$ with $g_{11}(x)\mid x^{m}-1$ and $g(x)=\gcd(g_{11}(x),g_{12}(x))$, then $\dim C=m-\deg g(x)$. With the choice of $g_{22}(x)$ described in Lemma 2.4, we see that the dimension of $C$ in Theorem 2.2 is consistent with the result by Séguin.
@@ 3 Euclidean LCD quasi-cyclic codes of index 2
Let $C$ be a quasi-cyclic code of index $2$. Let $(g_{11}(x),g_{12}(x))$ and $(0,g_{22}(x))$ be the generators of $C$ satisfying Conditions $(\ast)$. Then $C$ is Euclidean LCD if and only if all of the following conditions are true:
@@
Assume that $l$ is not self-reciprocal. This implies that there exists $h_{j}$ such that $h_{j}\mid l$ and $h_{j}^{*}\nmid l$. Since $h_{j}\mid l$, it follows that $h_{j}\nmid\text{lcm}(g_{11},g_{22})$. In particular, $h_{j}\nmid g_{11}$ and $h_{j}\nmid g_{22}$. Then $g_{11}(\xi^{v_{j}})\neq 0$, $g_{22}(\xi^{v_{j}})\neq 0$ and $\operatorname{rank}(G_{j}^{\prime})=2$. This implies that $C_{j}^{\prime}=(H_{j}^{\prime})^{2}.$
@@ 1. Since $g$ and $l$ are self-reciprocal,
Furthermore, the reciprocal of $x^{m}-1$ is $(x^{m}-1)$, so we can rewrite
@@
2. $h_{j}\mid l$. This condition implies that $h_{j}\nmid g$ and $h_{j}^{*}\nmid g$. Then $C_{j}^{\prime}$ and $C_{j}^{\prime\prime}$ are 2-dimensional, which implies that $C_{j}^{\prime\perp_{e}}$ and $C_{j}^{\prime\prime\perp_{e}}$ are 0-dimensional. Then $C_{j}^{\prime}\cap C_{j}^{\prime\prime\perp_{e}}=\{\mathbf{0}\}$ and $C_{j}^{\prime\prime}\cap C_{j}^{\prime\perp_{e}}=\{\mathbf{0}\}$.
@@
In [8], the authors showed that any linear code is equivalent to a Euclidean LCD code for $q>3$. Using Theorem 3.1, we construct several binary ($q=2$) (in Table 1) and ternary $(q=3)$ quasi-cyclic Euclidean LCD codes.
$C$ $g_{11}(x)$ $g_{12}(x)$ $g_{22}(x)$ Remark
$[46,23,10]$ $x+1$ $1+x^{2}+x^{6}+x^{8}$ $x^{9}+x^{13}+x^{15} +$$x^{20}$ $(x^{11}+x^{10}+x^{6}+x^{5}$ $x^{4}+x^{2}+1)$$(x^{11}$ $x^{9}+x^{7}+x^{6} +$$x^{5}+x+1)$ Near-optimal
$[50,24,10]$ $(x+1)(x^{4}+x^{3}$ $x^{2}+x+1)$ $(x+1)(1+x^{5}+$ $x^{7}+x^{14})$ $(x+1)(x^{20}+x^{15}$ $x^{10}+x^{5}+1) -$
$[50,28,8]$ $x+1$ $(x+1)(1+x^{5}$ $x^{7}+x^{8})$ $(x+1)(x^{20}+x^{15}$ $x^{10}+x^{5}+1)$ BKLC
$[54,21,12]$ $(x+1)$ $(x^{6}+x^{3}+1)$ $(x+1)$ $(x^{6}+x^{3}+1)$ $(x^{2}+x+1)$ $(x^{4}+x+1)$ $(x^{2}+x+1)$ $(x^{6}+x^{3}+1)$ $(x^{18}+x^{9}+1)$ -
$[54,25,11]$ $(x+1)$ $(x^{2}+x+1)$ $(x^{2}+x+1)$ $(x^{13}+x^{9}+x^{7}$ $x^{6}+x^{5}+1)$ $(x^{2}+x+1)$ $(x^{6}+x^{3}+1)$ $(x^{18}+x^{9}+1)$ New
$[54,26,8]$ $(x+1)$ $(x^{2}+x+1)$ $(x+1)$ $(x^{8}+x^{7}+x^{5}+1)$ $(x+1)$ $(x^{6}+x^{3}+1)$ $(x^{18}+x^{9}+1)$ New
Table 1: Binary quasi-cyclic (Euclidean) LCD codes, where BKLC means that the minimum distance of the code is equal to the minimum distance of Best Known Linear Code according to the Grassl’s table, “New” means codes are new to the codes given in [14], [24], [44] with given parameters. “Near-optimal” means that the minimum distance is one less than the optimal distance for linear codes (it might be optimal as an LCD code), and it is better than the LCD code [46,23,9] in [44].
Example 1.
Let $q=3$ and $m=13$. Then $x^{13}-1=(x+2)(x^{3}+2x+2)(x^{3}+x^{2}+2)(x^{3}+x^{2}+x+2)(x^{3}+2x^{2}+2x+2)$. Suppose $C$ is generated by
$\begin{split}g_{11}(x)&=x^{3}+2x+2,\\ g_{12}(x)&=x^{7}+2x^{4}+x+2,\\ g_{22}(x)&=(x^{3}+x^{2}+2)(x^{3}+x^{2}+x+2)(x^{3}+2x^{2}+2x+2).\end{split}$
Then, by Theorem 3.1, $C$ is a quasi-cyclic ternary LCD code with parameters $[26,14,7]$. Its minimal distance is the same as that of the Best Known Linear Code (BKLC) as per Grassl’s Table. The dual code has parameters $[26,12,8]$.
Example 2.
Let $q=3$ and $m=13$. Suppose $C$ is generated by
$\begin{split}g_{11}(x)&=x+2,\\ g_{12}(x)&=x^{5}+2x^{4}+2x^{3}+x^{2}+x+1,\\ g_{22}(x)&=(x^{3}+x^{2}+x+2)(x^{3}+2x^{2}+2x+2).\end{split}$
Then, by Theorem 3.1, $C$ is a quasi-cyclic ternary LCD code with parameters $[26,19,4]$. This is a near optimal code as per Grassl’s Table.
@@ 4 One-generator quasi-cyclic codes of index 2
In view of Lemma 2.4, we can assume that $C$ is generated by two elements $(g_{11}(x),\tilde{g}_{12}(x))$ and $(0,g_{22}(x))$ satisfying Conditions $(\ast)$, where $\tilde{g}_{12}(x)=g_{12}(x)\mod{g_{22}(x)}$.
@@ 2. The conditions $\gcd(g_{11}^{\prime},g_{12})=1$ and $\gcd(g_{11}^{\prime},\bar{g}_{12})=1$ imply that
We recall from Lemma 2.4 we can assume that the code $C$ is generated by two elements $(g_{11}(x),\tilde{g}_{12}(x))$ and $(0,g_{22}(x))$ satisfying Conditions $(\ast)$, where
@@ 5 Symplectic LCD quasi-cyclic codes of index 2
Assume that all the notations are the same as in the previous sections. First, we recall some definitions and results, for details see [11, Section V]. For $x=(x_{1}|x_{2})$ and $y=(y_{1}|y_{2})$ in $F^{2m}$, where $x_{i},y_{i}\in F^{m}$ for $i=1,2$, we have
@@
Let $C$ be a QC code of length $2m$ and index $2$ with CRT decomposition given in (2). By extending the map $\tau$ canonically to the vector spaces $(F_{i})^{2},(H_{j}^{\prime})^{2},(H_{j}^{\prime\prime})^{2}$ and applying the maps component-wise to (3), we obtain symplectic dual of $C$ (see [11, Proposition V.1]).
@@
Using the above characterization, we have the following characterization of symplectic LCD QC codes in terms of constituents (see [11, Eq. V.7]).
@@
Let $C$ be a quasi-cyclic code of index $2$. Let $(g_{11}(x),g_{12}(x))$ and $(0,g_{22}(x))$ be the generators of $C$ satisfying Conditions $(\ast)$. Then $C$ is symplectic LCD if and only if all of the following conditions are true:
@@
Furthermore, the reciprocal of $x^{m}-1$ is $(x^{m}-1)$, so we can rewrite
@@
Example 3.
Let $m=17$, $q=2$. Then $x^{m}-1=(x+1)(x^{8}+x^{5}+x^{4}+x^{3}+1)(x^{8}+x^{7}+x^{6}+x^{4}+x^{2}+x+1)\in\mathbb{F}_{2}[x]$. Take
$\begin{split}g_{11}=&x+1,\\ g_{22}=&(x+1)(x^{8}+x^{7}+x^{6}+x^{4}+x^{2}+x+1),\\ g_{12}=&(x+1)(x^{4}+x^{3}+x^{2}+x+1).\end{split}$
Then the code generated by $(g_{11},g_{12})$ and $(0,g_{22})$ is a symplectic LCD code with parameters $[34,24,4]$. It is optimal as per Grassl’s table.
@@
Restricting to the case $r_{11}(x)=r_{22}(x)=1$ in Theorem 5.3, we can get the following sufficient condition for a code to be symplectic LCD.
Corollary 5.7.
Let $C$ be a quasi-cyclic code of index $2$. Let $(g_{11}(x),g_{12}(x))$ and $(0,g_{22}(x))$ be the generators of $C$ satisfying Conditions $(\ast)$. If the polynomial $g(x)=\gcd(g_{11}(x),g_{22}(x))=\gcd(g_{11}(x),g_{11}^{*}(x))=\gcd(g_{22}(x),g_{22}^{*}(x))$ is self-reciprocal and $g_{11}(x)g_{22}(x)$ is self-reciprocal, then $C$ is symplectic LCD.
In the following example, we construct a quasi-cyclic (symplectic) LCD code using Corollary 5.7.
Example 4.
Let $m=15$, $q=2$. Then $x^{m}-1=(x+1)(x^{2}+x+1)(x^{4}+x+1)(x^{4}+x^{3}+1)(x^{4}+x^{3}+x^{2}+x+1)\in\mathbb{F}_{2}[x]$. Take
$\begin{split}g_{11}=&(x^{4}+x^{3}+x^{2}+x+1)(x^{4}+x+1),\\ g_{22}=&(x^{4}+x^{3}+x^{2}+x+1)(x^{4}+x^{3}+1),\\ g_{12}=&x^{4}+x^{3}+x^{2}+x+1.\end{split}$
Then the code generated by $(g_{11},g_{12})$ and $(0,g_{22})$ is a symplectic LCD code with parameters $[30,14,5]$.